A is directly proportional to B. When A is 7, B is 2. What is B when A is 10?
P is inversely proportional to Q. P is 5 when Q is (1/10). What is P when Q is 0.01?
I am confused? Do you set this up as a cross multiplication? Please help. Thankyou Bill

Answer by stanbon(75887)

A is directly proportional to B.
A = kB
----------
When A is 7, B is 2.
Use this information to solve for "k":
7 = k*2
k = 7/2
----------------
Therefore the equations that relates A and B is A = (7/2)B
What is B when A is 10?
10 = (7/2)B
B = 20/7
===========================
Next Problem:
P is inversely proportional to Q
P = k/Q
--------------
P is 5 when Q is (1/10)
5 = k/(1/10)
k = 1/2
--------
EQUATION:
P = 1/(2Q)
What is P when Q is 0.01?
P = 1/(0.02)
P = 50
============
Cheers,
Stan H.